OpenStudy (anonymous):

PLEASEE HELP ILL GIVE MEDALSS!!!!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (ibbutibbu.):

-2 = (3)4 + b -2 = 12 + b -14 = b so therefore the equation is y = 3x -14

OpenStudy (ibbutibbu.):

or 3x - y = -14

OpenStudy (anonymous):

how do I give u a medal?? nd thank you <3

OpenStudy (anonymous):

if both of these linnes are paralell then thier slope is same...

OpenStudy (anonymous):

just clik the best response button in coment..